OpenStudy (anonymous):

Find the dilated function of y=2x^2 -5 with the scale factor of 3.

OpenStudy (anonymous):

A. y=2x^2 +15 B. y=6x^2 -15 C. y=6x^2 -5 D. y=2x^2 -15

OpenStudy (anonymous):

C

OpenStudy (anonymous):

it was B

OpenStudy (anonymous):

unless you had an x after that -5, it shouldn't have been

myininaya (myininaya):

In the future, it would be better if answers weren't just given. Please guide the asker in the future.